# Robust Online Learning

### v1

@@ 2 Setup and Problem Formulation
Consider a sequence $S=X_{1},Y_{1},\cdots,X_{T},Y_{T}$ for $T\leq\infty$. For any finite $t\leq T$, we denote a prefix of $S$ by $S_{<t}=X_{1},Y_{1},\cdots,X_{t-1},Y_{t-1}$. We respectively define the mistake bound and the regret of a learner $\mathcal{A}$ with respect to $\mathcal{H}$ on a sequence $S$ as follows:
@@
$\mathrm{R}_{\mathcal{H}}(\mathcal{A},S)=\sum_{t=1}^{T}\mathbf{1}[\mathcal{A}(S_{<t})(Z_{t})\neq Y_{t}]-\inf_{h\in\mathcal{H}}\sum_{t=1}^{T}\mathbf{1}[h(Z_{t})\neq Y_{t}]$ (3)
@@ 3.1 Orientation Game
Figure 1: A $\mathcal{U}$-adversarial tree of depth 2. For each $k\leq 2$ and $\mathbf{u}\in\{0,1\}^{k}$ we have $\mathcal{U}(x_{\mathbf{u}}^{0})\cap\mathcal{U}(x_{\mathbf{u}}^{1})\neq\emptyset$. The tree is shattered $\mathcal{H}$ if each of its root-to-leaf paths are realizable by $\mathcal{H}$. For example, there must exist $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}h_{01}$$\in\mathcal{H}$ such that ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}h_{01}}(z)={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0}$ for all $z\in\mathcal{U}(x_{\emptyset}^{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0}})$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}h_{01}}(z)={\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}1}$ for all $z\in\mathcal{U}(x_{0}^{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}1}})$.
@@ 5 Regret bounds for agnostic robust online learning
Algorithm 3 Expert $\hat{Y}_{t}(i_{1},\cdots,i_{L})$
Input: Hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$, perturbation set $\mathcal{U}$, orientation learner $\eta:(\mathcal{X}^{2}_{\mathcal{U}}\times\mathcal{Y})^{*}\times\mathcal{X}^{2}\to\{0,1\}$, indices $i_{1},\cdots,i_{L}\in[T]$
for $t\geq 1$ do
   Receive $Z_{t}$ for $y\in\mathcal{Y}$ do
   end for
    $\tilde{Y}_{t}\leftarrow y$
   end if
 else
    $\tilde{Y}_{t}\leftarrow 1$
   end if
 if $t\in\{i_{1},\cdots,i_{L}\}$ then
    $\hat{Y}_{t}\leftarrow 1-\tilde{Y}_{t}$
   end if
 else
    $\hat{Y}_{t}\leftarrow\tilde{Y}_{t}$
   end if
  Receive $X_{t}$if $t\in\{i_{1},\cdots,i_{L}\}$ then
      Let $X_{\tau}^{\hat{Y}_{t}}\leftarrow X_{t}$  Let $X_{\tau}^{1-\hat{Y}_{t}}$ be an arbitrary $x\in P_{1-\hat{Y}_{t}}$ such that $f(X_{\tau}^{0},X_{\tau}^{1})=1-\hat{Y}_{t}$  Let $\xi_{\tau}\leftarrow((X_{\tau}^{0},X_{\tau}^{1}),\hat{Y}_{t}$)  Update $f(.,.)\leftarrow\eta(\xi_{1},\dots,\xi_{\tau},(.,.))$  Update $\tau\leftarrow\tau+1$
   end if
 $\mathcal{V}\leftarrow\mathcal{V}_{X_{t},\hat{Y}_{t}}^{\mathcal{U}}$
end for
So far, the adversary was restricted to choose realizable sequences. Here we consider a problem where the adversary does not have any restriction on its sequence. In this case, we are interested in finding an algorithm whose regret w.r.t. the best hypothesis in the class is sublinear. For this problem, we must allow the learner to have some private randomization by the same arguments made for classic online learning [17]. Then we have the following result on the achievable expected regret (Theorem 15). The proof of this result is by the general technique of agnostic to realizable reduction [3]. The idea of this technique is to create a set of “experts" such that for each input sequence and for each hypothesis in the class, there is an expert that predicts exactly the same as the hypothesis. Thus, it is enough to achieve small regret w.r.t. the best expert, which can be done using the algorithm of “prediction with expert advice" [4, Chapter 2]. Here we only focus on binary classes, however, we believe similar results can be proven by adapting the methods used for the non-robust online learning for multiclass classes [7, 9].
@@
Consider any hypothesis class $\mathcal{H}$ with $\mathrm{L}_{\mathcal{U}}(\mathcal{H})<\infty$. For any $h\in\mathcal{H}$, $\tau\leq T$, $X_{1},\cdots,X_{\tau}\in\mathcal{X}$, and for any $Z_{1},\cdots,Z_{\tau}$, there exists $L$ and $i_{1},\cdots,i_{L}\in\mathbb{N}$, such that the expert $Y_{t}(i_{1},\cdots,i_{L})$ defined in Algorithm 3 predicts the same as $h$ on $Z_{t}$ for all $t\leq\tau$.
@@
Assume we run the realizable online learner $\hat{Y}_{t}$ on the sequence $Z_{1},\cdots Z_{\tau}$ with the feedback
$X_{1},h(Z_{1}),\cdots,X_{\tau},h(Z_{\tau})$. We let $L$ be the number of mistakes made by the learner. By Theorem 10, $L\leq\mathrm{L}_{\mathcal{U}}(\mathcal{H})$. Furthermore, we define $i_{1},\cdots,i_{L}$ to be the rounds at which the learner makes a mistake. Now when we run $\hat{Y}_{t}(i_{1},\cdots,i_{L})$ on $Z_{1},\cdots,Z_{t}$, we will always predict according to $h$ because on rounds that the realizable learner makes a mistake we are flipping its prediction. Note that by keeping $\hat{Y}_{t}=Y_{t}=h(Z_{t})$, the update rules (e.g., the version space) in $\hat{Y}_{t}(i_{1},\cdots,i_{L})$ is the same as the realizable learner (a more formal proof is by a simple induction). ∎
Theorem 15.
For any binary hypothesis class $\mathcal{H}$ with finite $\mathrm{L}_{\mathcal{U}}(\mathcal{H})$, there exists a learning algorithm that achieves expected regret $\mathcal{O}(\sqrt{\mathrm{L}_{\mathcal{U}}(\mathcal{H})T\log(T)})$ in the agnostic robust online learning problem.
Proof.
Consider all possible experts for $L\in\{0,\cdots,\mathrm{L}_{\mathcal{U}}(\mathcal{H})\}$ and all indices $i_{1},\cdots,i_{L}\in[T]$. Thus, the total number of experts is $N=\sum_{L=0}^{\mathrm{L}_{\mathcal{U}}(\mathcal{H})}\binom{T}{L}\leq T^{\mathrm{L}_{\mathcal{U}}}$. By Lemma 14, for any sequence chosen by the adversary, for any $h\in\mathcal{H}$, there exists an expert whose behavior, and thus whose number of mistakes, is the same as $h$. This is specifically true for the hypothesis that achieves the optimal number of mistakes on the sequence. Thus, the regret of a learner w.r.t. the best expert is larger than its regret w.r.t. the best hypothesis in the class. Thus, it suffices to bound the former. To achieve this goal we can use the algorithm for the problem of learning with expert advice which achieves the expected regret of $\mathcal{O}(\sqrt{\log(N)T})=\mathcal{O}(\sqrt{\mathrm{L}_{\mathcal{U}}(\mathcal{H})T\log(T)})$ (see Appendix A for a short introduction to prediction with expert advice). ∎
We also have the following lower bound on the regret of any learner. We omit the proof here as it is a simple generalization of the proof for the classic agnostic online learning [3, Lemma 14].
Theorem 16.
For any binary hypothesis class $\mathcal{H}$, no learner can achieve (expected) regret smaller than $\Omega(\sqrt{\mathrm{L}_{\mathcal{U}}(\mathcal{H})T})$ in the agnostic robust online learning.
@@ 6 Robust online learning with uncertain perturbation sets
In previous sections we assumed the learner knows $\mathcal{U}$. In this section we weaken this assumption so that the learner does not fully know $\mathcal{U}$ but has some prior knowledge about it. Formally, we assume the true perturbation function $\mathcal{U}^{*}$ belongs to some finite $\mathcal{G}\subseteq(2^{\mathcal{X}})^{\mathcal{X}}$, which is known to the learner. Here we study this scenario for both realizable and agnostic learning of binary classes in finite number of rounds $T$. The overall idea is to run the algorithms developed in the previous sections for all possible perturbation functions in $\mathcal{G}$ and then consider them as experts that we can use to make predictions. We are now ready to state our results.
@@
We can also prove regret bounds in the agnostic setting similar to Theorem 17.
Theorem 19.
Any class $\mathcal{H}$ with $L^{*}:=\max_{\mathcal{U}\in\mathcal{G}}\mathrm{L}_{\mathcal{U}}(\mathcal{H})<\infty$ can be agnostically $\mathcal{U}^{*}$-robustly learned with expected regret bounded by $\mathcal{O}(\sqrt{T(L^{*}\log(T)+\log(|\mathcal{G}|))})$.
Proof.
For each $\mathcal{U}\in\mathcal{G}$ and for each $L\leq L^{*}$ and $i_{1},\cdots,i_{L}\leq L$, define expert $\mathcal{A}_{\mathcal{U}}(i_{1},\cdots,i_{L})$ that predicts according to the strategy $\hat{Y}_{t}(i_{1},\cdots,i_{L})$ assuming the perturbation function is $\mathcal{U}$. Again, if the version space become empty the expert will predict 1. By Lemma 14, we know for input sequence and for any $h\in\mathcal{H}$, there exists $L\leq\mathrm{L}_{\mathcal{U}^{*}}(\mathcal{H})\leq L^{*}$ and $i_{1},\cdots,i_{L}$ such that $\mathcal{A}_{\mathcal{U}^{*}}(i_{1},\cdots,i_{L})$ predicts the same as $h$ on that sequence. Thus, the regret of any learner with respect to these experts is bigger than the regret w.r.t. the best hypothesis in the class. Finally, the results follows by Theorem 20 with $N\leq T^{L}|\mathcal{G}|$. ∎

### v2

# arXiv:2602.06775v2[cs.LG] 26 Feb 2026

@@ 2 Setup and Problem Formulation
Consider a sequence $Z_{1},X_{1},Y_{1},\cdots,Z_{T},X_{T},Y_{T}$ for $T\leq\infty$. For any finite $t\leq T$, define $S_{<t}=Z_{1},X_{1},Y_{1},\cdots,Z_{t-1},X_{t-1},Y_{t-1}$. We respectively define the mistake bound and the regret of a learner $\mathcal{A}$ with respect to $\mathcal{H}$ on a sequence $S$ as follows:
@@
$R_{\mathcal{H}}(\mathcal{A},S)=\sum_{t=1}^{T}\mathbf{1}[\mathcal{A}(S_{<t})(Z_{t})\neq Y_{t}]-\min_{h\in\mathcal{H}}\sum_{t=1}^{T}l_{\mathcal{U}}(h,(X_{t},Y_{t}))$ (3)
@@ 3.1 Orientation Game
Figure 1: A $\mathcal{U}$-adversarial tree of depth 2. For each $k\leq 2$ and $\mathbf{u}\in\{0,1\}^{k}$ we have $\mathcal{U}(x_{\mathbf{u}}^{0})\cap\mathcal{U}(x_{\mathbf{u}}^{1})\neq\emptyset$. The tree is shattered by $\mathcal{H}$ if each of its root-to-leaf paths are realizable by $\mathcal{H}$. For example, there must exist $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}h_{01}$$\in\mathcal{H}$ such that ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}h_{01}}(z)={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0}$ for all $z\in\mathcal{U}(x_{\emptyset}^{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0}})$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}h_{01}}(z)={\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}1}$ for all $z\in\mathcal{U}(x_{0}^{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}1}})$.
@@ 5 Regret bounds for agnostic robust online learning
In the agnostic setting, we do not assume the sequence is realizable by the class $\mathcal{H}$. However, instead of minimizing the number of mistakes, we minimize the regret of the algorithm (Equation 3). Moreover, we allow the learner to be randomized, in which case we are interested in the expected regret.
To derive regret upper bounds in the agnostic setting, we use the technique of [9] where we compress the input sequence to the maximally realizable subsequence of the input sequence.
Theorem 14.
The optimal expected regret on any sequence $S=Z_{1},X_{1},Y_{1},\cdots,Z_{T},X_{T},Y_{T}$ is at most $\mathcal{O}(\sqrt{T\mathrm{L}_{\mathcal{U}}(\mathcal{H})\log(T)})$.
Proof.
For any $J\subseteq[T]$, let $S_{J}:=\{Z_{t},X_{t},Y_{t}:t\in J\}$ be the subsequence of $S$ on indices in $J$. Let $\mathcal{A}(z;S_{J})$ be the realizable learner that has only been trained using the subsequence $S_{J}$ (since the sequence might not be realizable, we assume the learner predicts 0 if the version space becomes empty). Here we assume the realizable learner is lazy and only updates its history (version space) when it makes a mistake, which is enough for achieving the same upper bound in the realizable setting. This is because each mistake still reduces the dimension of the version space by at least one, so the number of mistakes will still be bounded by the dimension of the hypothesis class. Let $h^{*}\in\mathcal{H}$ be the hypothesis that achieves $\min_{h\in\mathcal{H}}\sum_{t=1}^{T}l_{\mathcal{U}}(h,(X_{t},Y_{t}))$. Let $R=\{t\in[T]:l_{\mathcal{U}}(h^{*},(X_{t},Y_{t}))=0\}$. Then $S_{R}$ is realizable by $h^{*}$ and thus $\mathcal{H}$. Assume we play the robust online learning game on $S_{R}$, then by the guarantees of the (lazy) realizable learner, it makes $L\leq\mathrm{L}_{\mathcal{U}}(\mathcal{H})$ mistakes. Let $J$ be the rounds on which the lazy learner makes a mistake. Consider the prediction strategy on the whole sequence that at each round $t$ predicts $\mathcal{A}_{J}(Z_{t})=\mathcal{A}(Z_{t};S_{J_{<t}})$ with $J_{<t}=\{j\in J:j<t\}$. Notice that this means on rounds in $\R[T]\backslash$ R, we do not perform any updates on the learner, and on $R$ we only make updates when we make a mistake. Then by this definition, the mistakes on $S_{R}$ are exactly the subset $J$. Thus,
[…]
$\leq L+T-|R|$
$\leq\mathrm{L}_{\mathcal{U}}(\mathcal{H})+\sum_{t=1}^{T}l_{\mathcal{U}}(h^{*},(X_{t},Y_{t}))$
However, we do not know the sequence and thus $h^{*},R$, and $J$ in advance, so we need to consider all possibilities of $J$, that is all subsets of size at most $\mathrm{L}_{\mathcal{U}}(\mathcal{H})$ of $[T]$ whose count is at most $T^{\mathrm{L}_{\mathcal{U}}(\mathcal{H})}$ for large enough $T$. As our final algorithm, $\mathcal{A}_{agn}$, we run the prediction with expert advice algorithm of Theorem 18 over $\{A_{J}:J\subseteq T,|J|\leq\mathrm{L}_{\mathcal{U}}(\mathcal{H})\}$, which gives us the following guarantee:
$\sum_{t=1}^{T}\mathbb{E}[1[\mathcal{A}_{agn}(Z_{t})\neq Y_{t}]]$ $\leq\min_{J\in\binom{[T]}{\leq\mathrm{L}_{\mathcal{U}}(\mathcal{H})}}1[\mathcal{A}_{J}(Z_{t})\neq Y_{t}]+\mathcal{O}(\sqrt{T\log(T^{\mathrm{L}_{\mathcal{U}}(\mathcal{H})})})$
$\leq\mathrm{L}_{\mathcal{U}}(\mathcal{H})+\sum_{t=1}^{T}l_{\mathcal{U}}(h^{*},(X_{t},Y_{t}))+\mathcal{O}(\sqrt{T\mathrm{L}_{\mathcal{U}}(\mathcal{H})\log(T)})$
Thus, the expected regret is bounded by $\mathcal{O}(\sqrt{T\mathrm{L}_{\mathcal{U}}(\mathcal{H})\log(T)})$. ∎
@@
For any class $\mathcal{H}$ and any agnostic learner, there exists a sequence on which the optimal expected regret is at least $\Omega(\sqrt{T\mathrm{L}_{\mathcal{U}}(\mathcal{H})})$.
@@
The proof follows similar ideas to those of the proof for classical setting by [3] (Lemma 14). ∎
@@ 6 Robust online learning with uncertain perturbation sets
In previous sections we assumed the learner knows $\mathcal{U}$. In this section we weaken this assumption so that the learner does not fully know $\mathcal{U}$ but has some prior knowledge about it. Formally, we assume the true perturbation function $\mathcal{U}^{*}$ belongs to some finite $\mathcal{G}\subseteq(2^{\mathcal{X}})^{\mathcal{X}}$, which is known to the learner. Here we study this scenario for realizable learning in finite number of rounds $T$. The overall idea is to run the algorithms developed in the previous sections for all possible perturbation functions in $\mathcal{G}$ and then consider them as experts that we can use to make predictions. We are now ready to state our results.